18 August 2015: Tuesday

LDAP:

  • LDAP Update:
    • Two graduate student employees with missing info.  Found the matches and merged them.
    • Still have the one different name left from yesterday.
  • ACCOUNT Update:
    • No issues
  • ACL updates for Brian.

Backups:

  • Save Groups:
    • The usual failure.
  • Clones:
    • CoM/IS and Disk clones finished.
    • Tape clones are almost done.

VPN Change:

  • LDAP update updated – on test system.  Will let it run for a while before I throw it into production.
  • One-off massive update – oops – just dumping them out and running “common” is not going to work.  I will need to actually write a process.

Tags: ,

Leave a Reply

You must be logged in to post a comment.